OpenStudy (anonymous):

2. What is the simplified form of (3x + 2)(4x – 3)?

OpenStudy (callisto):

(3x + 2)(4x – 3) = 3x(4x – 3) +2(4x – 3) = 12x^2 -9x +8x -6 = 12x^2 -x -6

OpenStudy (anonymous):

12x^2 - 9x + 8x - 6 12x^2 -x -6 www.teachon.tk

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome